OpenStudy (anonymous):

Which of these have a different pattern, and not the same pattern -4,-1,2,5 1,4,9,16 2.5,5,7.5,10

OpenStudy (anonymous):

how do you get from -4 to -1?

OpenStudy (anonymous):

what's one way?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

?

OpenStudy (anonymous):

+3

OpenStudy (anonymous):

good, now from -1 to 2?

OpenStudy (anonymous):

+3?

OpenStudy (anonymous):

excellent! and from 2 to 5?

OpenStudy (anonymous):

+3?

OpenStudy (anonymous):

I have to figure out which one of those have a different pattern one that is not the same like the one that isn't just +3 +3 +3,

OpenStudy (anonymous):

yep, so the first one just add 3 at each step. what about the next? from 1 to 4?

OpenStudy (anonymous):

+3?

OpenStudy (anonymous):

great. now from 4 to 9?

OpenStudy (anonymous):

+5

OpenStudy (anonymous):

right and from 9 to 16?

OpenStudy (anonymous):

+7?

OpenStudy (anonymous):

awesome! so we have 2 different patterns. the thrid should fit with one of those in some way. how do we get from 2.5 to 5?

OpenStudy (anonymous):

I'm sorry I am having trouble with this one

OpenStudy (anonymous):

try adding 2.5, what do you get?

OpenStudy (anonymous):

5!

OpenStudy (anonymous):

and now from 5 to 7.5?

OpenStudy (anonymous):

Sorry, dumb moment there

OpenStudy (anonymous):

now worries, I have them constantly!

OpenStudy (anonymous):

+2.5

OpenStudy (anonymous):

no not now

OpenStudy (anonymous):

Lol!

OpenStudy (anonymous):

yes! and to get from 7.5 to 10?

OpenStudy (anonymous):

+2.5!

OpenStudy (anonymous):

you're on it! so the first was +3 at each step and the thrid was +2.5 at each step and the second was +3 then +5 then +7. which is different?

OpenStudy (anonymous):

Could you help me with one more thing? Which of these represent the function of x that is equal to the number 3 subtracted from the sum of twice the value of x and 4? f(x)=3-(2x+4) F(X)=3+(2x-4) f(x)=(2x+4)-3 f(x)=(2x-4)+3 I just need some help on how I would figure it out

OpenStudy (anonymous):

what's being subtracted?

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

Yes, I am sorry I stepped away for a minute

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

so which choices have 3 being subtracted from something?

OpenStudy (anonymous):

the first and 3rd one

OpenStudy (anonymous):

nope, only the 3rd. in the first something is being subtracted from 3.

OpenStudy (anonymous):

Oh I'm sorry I didn't even realize that

OpenStudy (anonymous):

you should also notice that 3 is being subtracted from a sum (in the 3rd choice). The sum of 2 times x and 4. 2 times x is twice x.

OpenStudy (anonymous):

So my answer would be the 3rd one?

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I think it would be